OpenStudy (anonymous):

2/3 divided by 8/7 =

OpenStudy (anonymous):

Remember when we divide fractions, the second fraction we flip and multiply

OpenStudy (anonymous):

\[\huge \frac{ \frac{ 2 }{ 3 } }{ \frac{ 8 }{ 7 } } \implies \frac{ 2 }{ 3 } \times \frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

no i just multiply the two top one and then two bottom??

OpenStudy (anonymous):

Yeah multiplying fractions is easy, numerator by numerator, denominator by denominator :) so you're right!

OpenStudy (anonymous):

thank you so much